OpenStudy (anonymous):

simplify: sin(-x)cot(x) / sin((pi/2)-x)

OpenStudy (anonymous):

\[\frac{ \sin \left( -x \right)\cot x }{ \sin \left( \frac{ \pi }{ 2 } -x\right) }\] \[=\frac{ -\sin x \cot x }{ \cos x }=\frac{ -\sin x \cos x }{ \cos x \sin x }=?\]